Could you please help me. How can I solve this system of equation using the substitution method.
-5x+y=-3
3x-8y=24

Answer by jim_thompson5910(35256)

Let's solve for y in equation 1


-5x+y=-3


-5x+y-5x=-3-5x


y=-3-5x


----------------------------------------------------------------------------------------


Now we'll use y=-3-5x to plug into the other equation


3x-8y = 24


3x-8( y ) = 24


3x-8( -3-5x ) = 24 ... replace y with -3-5x; solve for x


3x-8(-3)+8(-5x) = 24


3x+24-40x = 24


3x-40x+24 = 24


-37x+24 = 24


-37x+24-24 = 24-24


-37x = 0


-37x/(-37) = 0/(-37)


x = 0


----------------------------------------------------------------------------------------


Now that we know x = 0, we can use this to find y


y=-3-5x


y=-3-5*0 ... replace x with 0


y=-3-0


y=-3


----------------------------------------------------------------------------------------


So together, x = 0 and y = -3.


They form the ordered pair (0,-3)


Visually, the two equations will graph 2 straight lines that meet at (0,-3)